1.The angle of refraction of a light ray is 1.4 times larger than the angle of incidence.What is the refractive index of the medium when i =20°


1
Expert's answer
2021-06-24T16:13:41-0400
"n=\frac{\sin 28}{\sin 20}=1.37"